. In a school 720 are boys and 40% are girls. Find the number of girls?

Answer by rfer(16322)

720/0.60=1200 total
1200*0.4=480 girls
-----------------------
720/0.6=x/0.4
0.6x=720*0.4
x=288/0.6
x=480 girls
You can work it two ways.